# Finite groups with some subgroups of prime power order satisfying the partial $\Pi$-property

### v1

@@ 1 Introduction
Chen and Guo proved in [4, Proposition 1.6] that a group $G$ in which every subgroup of order $p$, $p$ a prime, and every cyclic subgroup of order $4$ (if $p=2$ and $P$ is not quaternion-free), satisfy the partial $\Pi$-property in $G$ is $p$-supersoluble, that is, every chief factor of $G$ divisible by $p$ is cyclic.
Our first main result provides a structural information of a group $G$ in which every subgroup of order $p^{2}$ satisfies the partial $\Pi$-property. Without loss of generality, we may assume that the Sylow $p$-subgroups of $G$ have order at least $p^{2}$ and $O_{p^{\prime}}(G)=1$.
@@
Le $p$ be a prime, and let $P$ be a Sylow $p$-subgroup of a group $G$. Assume that $O_{p^{\prime}}(G)=1$ and $|P|\geq p^{2}$. If every subgroup of $P$ of order $p^{2}$ satisfies the partial $\Pi$-property in $G$, then $P=O_{p}(G)$ and $G=P\rtimes H$, the semidirect product of $P$ and $H$, where $H\in{\rm Hall}_{p^{\prime}}(G)$, and one of the following statements holds:
@@
Chen and Guo proved in [4, Proposition 1.4] that if the maximal subgroups of the Sylow $p$-subgroups of a group $G$ satisfy the $\Pi$-property in $G$, then either $G$ is $p$-supersoluble or the order of the Sylow $p$-subgroups of $G$ is at most $p$.
Our second main result describes the structure of a group $G$ in which the second maximal subgroups of the Sylow $p$-subgroups have the partial $\Pi$-property in $G$. We may assume that the Sylow $p$-subgroups of $G$ have order at least $p^{2}$ and $O_{p^{\prime}}(G)=1$.
@@
(2) If $G$ is $p$-soluble and $|P|=p^{2}$, then $P=O_{p}(G)$ is a minimal normal subgroup of $G$;
@@
Let $P\in{\rm Syl}_{p}(G)$ and let $d$ be a power of $p$ such that $1<d<|P|$. Assume that every subgroup of $P$ of order $d$ satisfies the partial $\Pi$-property in $G$, and assume further that every cyclic subgroup of $P$ of order $4$ satisfies the partial $\Pi$-property in $G$ when $d=2$ and $P$ is not quaternion-free. If the $p$-rank of $G$ is greater than $1$, then $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$. Furthermore, if $V$ is an irreducible $\mathbb{F}_{p}[H]$-submodule of $P/\Phi(P)$, and write
$dim_{\mathbb{F}_{p}}V=k$, $n=\log_{p}d-\log_{p}|\Phi(P)|$, and $m=\log_{p}|P/\Phi(P)|$,
@@ 2 Preliminaries
Assume that $H\cap G^{*}_{1}=1$. For every $G$-chief factor $G_{i}/G_{i-1}$ with $1<i\leq r$, we see that either $P=HG^{*}_{i-1}$ or $HG^{*}_{i-1}$ is a maximal subgroup of $P$. In both cases, $P$ normalises $HG^{*}_{i-1}$ and hence $P\leq N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})$. Since $|G:N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number, it follows that $HG^{*}_{i-1}\cap G^{*}_{i}$ is a normal subgroup of $G$. Therefore $H$ covers or avoids $G^{*}_{i}/G^{*}_{i-1}$. This implies that $H$ is a partial CAP-subgroup of $G$, as desired. Hence we can assume that $H\cap G^{*}_{1}=G^{*}_{1}$. By Lemma 2.1, the hypotheses are inherited by $G/G^{*}_{1}$. By induction, $H/G^{*}_{1}$ is is a partial CAP-subgroup of $G/G^{*}_{1}$. Hence $G/G^{*}_{1}$ has a chief series
@@
such that $H/G^{*}_{1}$ covers or avoids every $G/G^{*}_{1}$-chief factor $G_{j}/G^{*}_{1}\big{/}G_{j-1}/G^{*}_{1}$ of $\varGamma_{G/G^{*}_{1}}$, $2\leq j\leq n$. Therefore, $G_{j}/G^{*}_{1}\leq HG_{j-1}/G^{*}_{1}$ or $(H\cap G_{j})/G^{*}_{1}\leq G_{j-1}/G^{*}_{1}$. Consequently, $G_{j}\leq HG_{j-1}$ or $H\cap G_{j}\leq G_{j-1}$. This means that $H$ covers or avoids every $G$-chief factor of a chief series of $G$. Thus $H$ is a partial CAP-subgroup of $G$, as desired. ∎
Recall that a subgroup $H$ of $G$ is said to be complemented in a group $G$ if there exists a subgroup $K$ of $G$ such that $G=HK$ and $H\cap K=1$. In this case, $K$ is called a complement of $H$ in $G$.
@@
(1$\Rightarrow$(2). We argue by induction on $|G|$. If $H=1$ there is nothing to prove. By the Schur-Zassenhaus Theorem ([6, Theorem A.11.3]), $P$ is complemented in $G$ by a Hall $p^{\prime}$-subgroup $U$ of $G$. Hence we may assume that $H<P$. By Lemma 2.2, $G$ has a chief series
@@
passing through $P$ such that $|G:N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number for every $G$-chief factor $G^{*}_{i}/G^{*}_{i-1}$ of $\varOmega_{G}$, $1\leq i\leq n$. In particular, $|G:N_{G}(H\cap G^{*}_{1})|$ is a $p$-number, and so $G=PN_{G}(H\cap G^{*}_{1})$. Since $P$ is elementary abelian, we have $(H\cap G_{1}^{*})^{G}=H^{P}\cap G_{1}^{*}=H\cap G_{1}^{*}$. The minimality of $G_{1}^{*}$ implies that $H\cap G_{1}^{*}=G_{1}^{*}$ or $1$. If $H\cap G_{1}^{*}=G_{1}^{*}$, then $G_{1}^{*}\leq H$. By Lemma 2.1, $H/G_{1}^{*}$ satisfies the partial $\Pi$-property in $G/G_{1}^{*}$. By induction, $H/G_{1}^{*}$ has a complement in $G/G_{1}^{*}$, $K/G_{1}^{*}$ say. Thus $G=HK$ and $H\cap K=G_{1}^{*}$. Hence $P=H(P\cap K)$ and $P\cap K\unlhd G$. By Maschke’s theorem ([6, Theorem A.11.4]), there exists a normal subgroup $L$ of $G$ such that $P\cap K=G_{1}^{*}\times L$. In this case, $LU$ is a complement of $H$ in $G$, as required.
@@
(2$\Rightarrow$(1). Suppose that $H$ is complemented in $G$. We show that $H$ satisfies the $\Pi$-property in $G$ by induction on $|G|$. Let $B$ be a complement of $H$ in $G$. Then $P\cap B\unlhd G$ and $P=H\times(P\cap B)$. Since $H(P\cap B)/P\cap B$ is complemented in $G/P\cap B$, it follows that $H(P\cap B)/(P\cap B)$ satisfies the partial $\Pi$-property in $G/(P\cap B)$. Now, if $C/D$ is a $G$-chief factor below $P\cap B$, then $HD\cap C=D$ is normal in $G$. Consequently, $H$ satisfies the partial $\Pi$-property in $G$. ∎
@@
Let $H$ be a $p^{\prime}$-group and let $V$ be a faithful and irreducible $\mathbb{F}_{p}[H]$-module. Assume that $dim_{\mathbb{F}_{p}}V$ is prime. Then $H$ is cyclic if and only if $V$ is not absolutely irreducible.
@@ 3 Proofs
By Theorem 1.2, $G$ is $p$-soluble with $p$-length at most $1$. Since $O_{p^{\prime}}(G)=1$, it follows that $P$ is normal in $G$ and so the Schur-Zassenhaus Theorem implies that $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$. Since $G$ has a minimal normal subgroup of order $d$, it follows from Lemma 2.3 that every minimal normal subgroup of $G$ is elementary abelian of order $d$. Thus $N=\mathrm{Soc}(G)\leq P=O_{p}(G)$ and $N$ is a completely reducible $G$-module by Maschke’s theorem.
@@
Let $N$ be a minimal normal subgroup of $G$ contained in $\Phi(P)$. Clearly, $\frac{d}{|N|}\geq p$ since $|N|\leq\frac{d}{p}$. As $N\leq\Phi(P)\leq\Phi(G)$ and the $p$-rank of $G$ is greater than one, we can apply [6, Theorem IV.8.6] to conclude that the $p$-rank of $G/N$ is greater than $1$. By Lemma 2.1, every subgroup of $P/N$ of order $\frac{d}{|N|}$ satisfies the partial $\Pi$-property in $G/N$. If $p\geq 3$, then $G/N$ satisfies the hypotheses of the lemma. The minimal choice of $G$ implies that $\frac{d}{|N|}\geq p^{2}|\Phi(P/N)|=p^{2}|\Phi(P)/N|$, and hence $d\geq p^{2}|\Phi(P)|$, a contradiction. This yields that $p=2$ and $|N|=\frac{d}{2}$.
@@
If $G/N$ is $2$-supersoluble, then $G$ is $2$-supersoluble, a contradiction. Hence $G/N$ is not $2$-supersoluble. Let $M$ be a normal subgroup of $G$ of minimal order containing $N$ such that $M/N\nleq Z_{{\mathcal{U}}}(G/N)$ and $M\leq P$. Let $U/N$ be a minimal normal subgroup of $G/N$ contained in $M/N$. Then $U/N=M/N\cap Z_{{\mathcal{U}}}(G/N)$. By Lemma 2.4, $M/N$ possesses a cyclic subgroup $\langle x\rangle N/N$ of order $2$ or $4$ (when $M/N$ is not quaternion-free) such that $x\in M\setminus U$.
Denote with bars the images in $\overline{G}$. We will show that the exponent of $\overline{M}$ is $2$ or $4$ (when $\overline{M}$ is not quaternion-free). Let $\overline{D}$ be a Thompson critical subgroup of $\overline{M}$. If $\Omega(\overline{D})<\overline{M}$, then $\Omega(\overline{D})\leq\overline{U}$ by the minimal choice of $M$. Thus $\Omega(\overline{D})\leq Z_{{\mathcal{U}}}(\overline{G})$. By Lemma 2.5, we have that $\overline{M}\leq Z_{{\mathcal{U}}}(\overline{G})$, a contradiction. Hence $\overline{M}=\overline{D}=\Omega(\overline{D})$. If $\overline{M}$ is a non-abelian quaternion-free $2$-group, then $\overline{M}$ has a characteristic subgroup $\overline{M_{0}}$ of index $2$ by Lemma 2.7. The minimal choice of $M$ implies $M_{0}=U$. Thus $\overline{M}\leq Z_{{\mathcal{U}}}(\overline{G})$, which is impossible. Hence if $\overline{M}$ is a non-abelian $2$-group, then $\overline{M}$ is not quaternion-free. By Lemma 2.6, the exponent of $\overline{M}$ is $2$ or $4$ (when $\overline{M}$ is not quaternion-free), as desired.
@@
It is clear that $P$ is a completely reducible $\mathbb{F}_{p}[H]$-module by Maschke’s theorem. Write $P=V_{1}\times\cdots\times V_{t}$, where $V_{i}$ is an irreducible $\mathbb{F}_{p}[H]$-submodule of $P$ for all $1\leq i\leq t$. Write $\mathrm{dim}V_{i}=n_{i}$ and suppose that $n_{1}\leq n_{2}\leq\cdots\leq n_{t}$. By Lemma 2.3, we have $d\geq p^{n_{t}}$. Since $G$ is not $p$-supersoluble, it turns out that $d\geq p^{n_{t}}\geq p^{2}$ by [4, Proposition 1.6].
(1) We argue by induction on $|G|$. If $d=p^{n_{t}}$, then $n_{1}=n_{2}=\cdots=n_{t}$ by Lemma 2.3. In this case, the conclusion follows by Lemmas 2.12 and 2.13. We may assume that $d>p^{n_{t}}$ and $t>1$.
Suppose that $t=2$. By Lemma 2.1, every subgroup of $P/V_{1}$ of order $\frac{d}{|V_{1}|}$ satisfies the partial $\Pi$-property in $G/N$. Applying Lemma 2.3, we see that $|P/V_{1}|=p^{n_{2}}\leq\frac{d}{p^{n_{1}}}$. Consequently, $|P|=p^{n_{1}+n_{2}}\leq d<|P|$, a contradiction. Therefore $t\geq 3$. Note that every subgroup of $P/V_{j}$ of order $\frac{d}{|V_{j}|}$ satisfies the partial $\Pi$-property in $G/V_{j}$ and the $p$-rank of $G/V_{j}$ is equal to $n_{t}>1$ for $j=1,2$. Applying the inductive hypothesis to $G/V_{1}$ and $G/V_{2}$, we deduce that $V_{1},...,V_{t}$ are isomorphic $\mathbb{F}_{p}[H]$-modules of the same dimension $k>1$. Applying the inductive hypothesis to $G/V_{1}$, we have $k\big{|}\gcd(\log_{p}d-k,\log_{p}|P|-k)$. Hence $k\big{|}\gcd(\log_{p}d,\log_{p}|P|)$, as wanted.
@@
$U$ and $V_{m}=\{u^{m}\varphi(u)|u\in U\}$, where $m=0,1,...,p-1$.
@@
Applying Theorem 1.2, we have that $G$ is $p$-soluble of $p$-length at most $1$. Since $O_{p^{\prime}}(G)=1$, it follows that $P=O_{p}(G)$ and so $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$, by the Schur-Zassenhaus Theorem.
@@
Assume that $|P|\geq p^{2}$ and $G$ is not $p$-supersoluble. Suppose that $|P|=p^{2}$. If $G$ is $p$-soluble, then $P=O_{p}(G)$ is a minimal normal subgroup of $G$ and $G$ is of type (2).
Assume that $|P|\geq p^{3}$. If $\frac{|P|}{p^{2}}=2$, then $|P|=8$. Since $G$ is not $2$-supersoluble, we can apply Theorem 1.2 to conclude that $P\cong Q_{8}$. Hence $G$ is of type (3).

### v2

@@ 1 Introduction
Chen and Guo proved in [4, Proposition 1.6] that a group $G$ in which every subgroup of a Sylow $p$-subgroup $P$ of order $p$, $p$ a prime, and every cyclic subgroup of order $4$ (if $p=2$ and $P$ is not quaternion-free), satisfy the partial $\Pi$-property is $p$-supersoluble, that is, every chief factor of $G$ divisible by $p$ is cyclic.
Our first main result provides a structural information of a group $G$ in which every subgroup of a Sylow $p$-subgroup of order $p^{2}$ satisfies the partial $\Pi$-property. Without loss of generality, we may assume that the Sylow $p$-subgroups of $G$ have order at least $p^{2}$ and $O_{p^{\prime}}(G)=1$.
@@
Let $p$ be a prime, and let $P$ be a Sylow $p$-subgroup of a group $G$. Assume that $O_{p^{\prime}}(G)=1$ and $|P|\geq p^{2}$. If every subgroup of $P$ of order $p^{2}$ satisfies the partial $\Pi$-property in $G$, then $P=O_{p}(G)$ and $G=P\rtimes H$, the semidirect product of $P$ and $H$, where $H\in{\rm Hall}_{p^{\prime}}(G)$, and one of the following statements holds:
@@
Chen and Guo proved in [4, Proposition 1.4] that if the maximal subgroups of a Sylow $p$-subgroup of a group $G$ satisfy the $\Pi$-property in $G$, then either $G$ is $p$-supersoluble or the order of the Sylow $p$-subgroups of $G$ is $p$.
Our second main result describes the structure of a group $G$ in which the second maximal subgroups of a Sylow $p$-subgroup satisfy the partial $\Pi$-property in $G$. We may assume that the Sylow $p$-subgroups of $G$ have order at least $p^{2}$ and $O_{p^{\prime}}(G)=1$.
@@
(2) $|P|=p^{2}$ and $P=O_{p}(G)$ is a minimal normal subgroup of $G$;
(3) $|P|=p^{2}$ and $G$ is non-$p$-soluble;
@@
Let $P\in{\rm Syl}_{p}(G)$ and let $d$ be a power of $p$ such that $1<d<|P|$. Assume that every subgroup of $P$ of order $d$ satisfies the partial $\Pi$-property in $G$, and assume further that every cyclic subgroup of $P$ of order $4$ satisfies the partial $\Pi$-property in $G$ when $d=2$ and $P$ is not quaternion-free. If $O_{p^{\prime}}(G)=1$ and the $p$-rank of $G$ is greater than $1$, then $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$. Furthermore, if $V$ is an irreducible $\mathbb{F}_{p}[H]$-submodule of $P/\Phi(P)$, and write
$\mathrm{dim}_{\mathbb{F}_{p}}V=k$, $n=\log_{p}d-\log_{p}|\Phi(P)|$, and $m=\log_{p}|P/\Phi(P)|$,
@@ 2 Preliminaries
Assume that $H\cap G^{*}_{1}=1$. For every $G$-chief factor $G_{i}/G_{i-1}$ with $1<i\leq r$, we see that either $P=HG^{*}_{i-1}$ or $HG^{*}_{i-1}$ is a maximal subgroup of $P$. In both cases, $P$ normalises $HG^{*}_{i-1}$ and hence $P\leq N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})$. Since $|G:N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number, it follows that $HG^{*}_{i-1}\cap G^{*}_{i}$ is a normal subgroup of $G$. Therefore $H$ covers or avoids $G^{*}_{i}/G^{*}_{i-1}$ with $1<i\leq r$. Thus $H$ covers or avoids every $G$-chief factor of $\varOmega_{G}$. This implies that $H$ is a partial CAP-subgroup of $G$, as desired. Hence we can assume that $H\cap G^{*}_{1}=G^{*}_{1}$. By Lemma 2.1, the hypotheses are inherited by $G/G^{*}_{1}$. By induction, $H/G^{*}_{1}$ is is a partial CAP-subgroup of $G/G^{*}_{1}$. Hence $G/G^{*}_{1}$ has a chief series
@@
such that $H/G^{*}_{1}$ covers or avoids every $G/G^{*}_{1}$-chief factor $G_{j}/G^{*}_{1}\big{/}G_{j-1}/G^{*}_{1}$ of $\varGamma_{G/G^{*}_{1}}$, $2\leq j\leq n$. Therefore, $G_{j}/G^{*}_{1}\leq HG_{j-1}/G^{*}_{1}$ or $(H\cap G_{j})/G^{*}_{1}\leq G_{j-1}/G^{*}_{1}$. Consequently, $G_{j}\leq HG_{j-1}$ or $H\cap G_{j}\leq G_{j-1}$ with $2\leq j\leq n$. Hence $H$ covers or avoids every $G$-chief factor of a chief series of $G$. Thus $H$ is a partial CAP-subgroup of $G$, as desired. ∎
Recall that a subgroup $H$ of a group $G$ is said to be complemented in $G$ if there exists a subgroup $K$ of $G$ such that $G=HK$ and $H\cap K=1$. In this case, $K$ is called a complement of $H$ in $G$.
@@
(1$\Rightarrow$(2). We argue by induction on $|G|$. If $H=1$ there is nothing to prove. By the Schur-Zassenhaus Theorem ([6, Theorem A.11.3]), $P$ is complemented in $G$ by a Hall $p^{\prime}$-subgroup $U$ of $G$. Hence we may assume that $1<H<P$. By Lemma 2.2, $G$ has a chief series
@@
passing through $P$ such that $|G:N_{G}(HG^{*}_{i-1}\cap G^{*}_{i})|$ is a $p$-number for every $G$-chief factor $G^{*}_{i}/G^{*}_{i-1}$ of $\varOmega_{G}$, $1\leq i\leq n$. In particular, $|G:N_{G}(H\cap G^{*}_{1})|$ is a $p$-number, and so $G=N_{G}(H\cap G^{*}_{1})P$. Since $P$ is elementary abelian, we have $(H\cap G_{1}^{*})^{G}=H^{P}\cap G_{1}^{*}=H\cap G_{1}^{*}$. The minimality of $G_{1}^{*}$ implies that $H\cap G_{1}^{*}=G_{1}^{*}$ or $1$. If $H\cap G_{1}^{*}=G_{1}^{*}$, then $G_{1}^{*}\leq H$. By Lemma 2.1, $H/G_{1}^{*}$ satisfies the partial $\Pi$-property in $G/G_{1}^{*}$. By induction, $H/G_{1}^{*}$ has a complement in $G/G_{1}^{*}$, $K/G_{1}^{*}$ say. Thus $G=HK$ and $H\cap K=G_{1}^{*}$. Hence $P=H(P\cap K)$ and $P\cap K\unlhd G$. By Maschke’s theorem ([6, Theorem A.11.4]), there exists a normal subgroup $L$ of $G$ such that $P\cap K=G_{1}^{*}\times L$. Let $U$ be a Hall $p^{\prime}$-subgroup of $G$. In this case, $LU$ is a complement of $H$ in $G$, as required.
@@
(2$\Rightarrow$(1). Suppose that $H$ is complemented in $G$. We show that $H$ satisfies the partial $\Pi$-property in $G$ by induction on $|G|$. It is no loss to assume that $1<H<P$. Let $B$ be a complement of $H$ in $G$. Then $P\cap B\unlhd G$ and $P=H\times(P\cap B)$. Since $H(P\cap B)/P\cap B$ is complemented in $G/P\cap B$, it follows that $H(P\cap B)/(P\cap B)$ satisfies the partial $\Pi$-property in $G/(P\cap B)$. Now, if $C/D$ is a $G$-chief factor below $P\cap B$, then $H\cap C\leq D$. Consequently, $H$ satisfies the partial $\Pi$-property in $G$. ∎
@@
Let $H$ be a $p^{\prime}$-group and let $V$ be a faithful and irreducible $\mathbb{F}_{p}[H]$-module. Assume that $\mathrm{dim}_{\mathbb{F}_{p}}V$ is prime. Then $H$ is cyclic if and only if $V$ is not absolutely irreducible.
@@ 3 Proofs
By Theorem 1.2, $G$ is $p$-soluble with $p$-length at most $1$. Since $O_{p^{\prime}}(G)=1$, it follows that $P$ is normal in $G$ and so the Schur-Zassenhaus Theorem implies that $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$. Since $G$ has a minimal normal subgroup of order $d$, it follows from Lemma 2.3(2) that every minimal normal subgroup of $G$ is elementary abelian of order $d$. Thus $N=\mathrm{Soc}(G)\leq P=O_{p}(G)$ and $N$ is a completely reducible $G$-module by Maschke’s theorem.
@@
Let $N$ be a minimal normal subgroup of $G$ contained in $\Phi(P)$. Clearly, $\frac{d}{|N|}\geq p$ since $|N|\leq\frac{d}{p}$. Note that the class of all $p$-supersoluble groups is a saturated formation. As $N\leq\Phi(P)\leq\Phi(G)$ and the $p$-rank of $G$ is greater than one, we can conclude that the $p$-rank of $G/N$ is greater than $1$. By Lemma 2.1, every subgroup of $P/N$ of order $\frac{d}{|N|}$ satisfies the partial $\Pi$-property in $G/N$. If $p\geq 3$, then $G/N$ satisfies the hypotheses of the lemma. The minimal choice of $G$ implies that $\frac{d}{|N|}\geq p^{2}|\Phi(P/N)|=p^{2}|\Phi(P)/N|$, and hence $d\geq p^{2}|\Phi(P)|$, a contradiction. This yields that $p=2$ and $|N|=\frac{d}{2}$.
@@
If $G/N$ is $2$-supersoluble, then $G$ is $2$-supersoluble, a contradiction. Hence $G/N$ is not $2$-supersoluble. Let $M$ be a normal subgroup of $G$ of minimal order containing $N$ such that $M/N\nleq Z_{{\mathcal{U}}}(G/N)$ and $M\leq P$. Let $U/N$ be a minimal normal subgroup of $G/N$ contained in $M/N$. Then $U/N=M/N\cap Z_{{\mathcal{U}}}(G/N)$. By Lemma 2.4, $M/N$ possesses a cyclic subgroup $\langle x\rangle N/N$ of order $2$ or $4$ such that $x\in M\setminus U$.
Denote with bars the images in $\overline{G}=G/N$. We will show that the exponent of $\overline{M}$ is $2$ or $4$ (when $\overline{M}$ is not quaternion-free). Let $\overline{D}$ be a Thompson critical subgroup of $\overline{M}$. If $\Omega(\overline{D})<\overline{M}$, then $\Omega(\overline{D})\leq\overline{U}$ by the minimal choice of $M$. Thus $\Omega(\overline{D})\leq Z_{{\mathcal{U}}}(\overline{G})$. By Lemma 2.5, we have that $\overline{M}\leq Z_{{\mathcal{U}}}(\overline{G})$, a contradiction. Hence $\overline{M}=\overline{D}=\Omega(\overline{D})$. If $\overline{M}$ is a non-abelian quaternion-free $2$-group, then $\overline{M}$ has a characteristic subgroup $\overline{M_{0}}$ of index $2$ by Lemma 2.7. The minimal choice of $M$ implies $M_{0}=U$. Thus $\overline{M}\leq Z_{{\mathcal{U}}}(\overline{G})$, which is impossible. Hence if $\overline{M}$ is a non-abelian $2$-group, then $\overline{M}$ is not quaternion-free. By Lemma 2.6, the exponent of $\overline{M}$ is $2$ or $4$ (when $\overline{M}$ is not quaternion-free), as desired.
@@
It is clear that $P$ is a completely reducible $\mathbb{F}_{p}[H]$-module by Maschke’s theorem. Write $P=V_{1}\times\cdots\times V_{t}$, where $V_{i}$ is an irreducible $\mathbb{F}_{p}[H]$-submodule of $P$ for all $1\leq i\leq t$. Write $\mathrm{dim}V_{i}=n_{i}$ and suppose that $n_{1}\leq n_{2}\leq\cdots\leq n_{t}$. By Lemma 2.3(1), we have $d\geq p^{n_{t}}$. Since $G$ is not $p$-supersoluble, it turns out that $d\geq p^{n_{t}}\geq p^{2}$.
(1) We argue by induction on $|G|$. If $d=p^{n_{t}}$, then $n_{1}=n_{2}=\cdots=n_{t}$ by Lemma 2.3(2). In this case, the conclusion follows by Lemmas 2.12 and 2.13. We may assume that $d>p^{n_{t}}$ and $t>1$.
Suppose that $t=2$. By Lemma 2.1, every subgroup of $P/V_{1}$ of order $\frac{d}{|V_{1}|}$ satisfies the partial $\Pi$-property in $G/N$. Applying Lemma 2.3(1), we see that $|P/V_{1}|=p^{n_{2}}\leq\frac{d}{p^{n_{1}}}$. Consequently, $|P|=p^{n_{1}+n_{2}}\leq d<|P|$, a contradiction. Therefore $t\geq 3$. Note that every subgroup of $P/V_{j}$ of order $\frac{d}{|V_{j}|}$ satisfies the partial $\Pi$-property in $G/V_{j}$ and the $p$-rank of $G/V_{j}$ is equal to $n_{t}>1$ for $j=1,2$. Applying the inductive hypothesis to $G/V_{1}$ and $G/V_{2}$, we deduce that $V_{1},...,V_{t}$ are isomorphic $\mathbb{F}_{p}[H]$-modules of the same dimension $k>1$. Applying the inductive hypothesis to $G/V_{1}$, we have $k\big{|}\gcd(\log_{p}d-k,\log_{p}|P|-k)$. Hence $k\big{|}\gcd(\log_{p}d,\log_{p}|P|)$, as wanted.
@@
$U\;\text{and}\;V_{m}=\{u^{m}\varphi(u)|u\in U\},\;\text{where}\;m=0,1,...,p-1.$
@@
Applying [10, Theorem 1.3], we have that $G$ is $p$-soluble of $p$-length at most $1$. Since $O_{p^{\prime}}(G)=1$, it follows that $P=O_{p}(G)$ and so $G=P\rtimes H$, where $H\in\mathrm{Hall}_{p^{\prime}}(G)$, by the Schur-Zassenhaus Theorem.
@@
Assume that $|P|\geq p^{2}$ and $G$ is not $p$-supersoluble. Suppose that $|P|=p^{2}$. If $G$ is $p$-soluble, then $P=O_{p}(G)$ is a minimal normal subgroup of $G$ and $G$ is of type (2). If $G$ is non-$p$-soluble, then $G$ is of type (3).
Assume that $|P|\geq p^{3}$. If $\frac{|P|}{p^{2}}=2$, then $|P|=8$. Since $G$ is not $2$-supersoluble, we can apply [4, Proposition 1.6] to conclude that $P\cong Q_{8}$. Hence $G$ is of type (4).